# Quantitative estimates: How well does the discrete Fourier transform approximate the Fourier transform on $\mathbb{R}$?

### v1

@@ 2.2 Polynomial decay in frequency
Figure 2: Logarithmic plots of $E^{[n]}_{h}(f)$ for $f(x)=e^{-2\pi|x|}$. Points are computed numerically for $n=2^{l}$, $l=7,\ldots,23$. We choose $p=n^{\frac{1}{2}}$, $p=6n^{1/10}$, and $p=10$. Reference lines represent the theoretical slope $h^{\beta}$ with $\beta=\frac{3}{2}$ that leads to $n^{-\frac{3}{4}}$, $n^{-\frac{27}{20}}$, and $n^{-\frac{3}{2}}$, respectively. Points align well with theoretical predictions.
@@ 3 Approximation rates
In this section we derive general and explicit bounds on the approximation error $E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|$ in (5) by decomposing the error into a time and a frequency component that are estimated separately. These components arise from approximating $\hat{f}(\tfrac{k}{p})$ first by a Riemannian sum and then truncating it as in (3),
$\hat{f}\big{(}\tfrac{k}{p}\big{)}\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}\approx h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=\sqrt{hn}(\operatorname{\mathcal{F}}\!f_{h,n})(k)\,.$ (8)
The left-hand side is the sampled (continuous) Fourier transform, the right-hand side is its approximation by the discrete Fourier transform. The idea is to estimate the discretization error $\hat{f}\big{(}\tfrac{k}{p}\big{)}-h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ by the decay of $\hat{f}$, and the truncation $h\sum_{j\in\mathbb{Z}\setminus[n]}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ by the decay of $f$.
@@ 3.1 Error decomposition into time and frequency components
The continuous functions on the real line are denoted by $\mathcal{C}=\mathcal{C}(\mathbb{R})$. We consider the Wiener amalgam space $W(\mathcal{C},\ell^{1})$ that consists of all continuous functions $f$ such that
$\|f\|_{W(\mathcal{C},\ell^{1})}:=\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|$ (9)
is finite, cf. [9, 10, 12]. If $f\in W(\mathcal{C},\ell^{1})$, then the sampling operator $f\mapsto(f(hj))_{j\in\mathbb{Z}}\in\ell^{1}(\mathbb{Z})$ is bounded for all $h>0$, and consequently the $p$-periodization
@@
converges absolutely and uniformly for every $p>0$. For $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, the Poisson summation formula $\sum_{l\in\mathbb{Z}}f(x+pl)=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi i\frac{k}{p}x}$ holds pointwise for all $x\in\mathbb{R}$, cf. [14, Lemma 4].
@@
$h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}=h\sum_{j\in[n]}(\mathcal{P}\!_{p}f)(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=(\mathcal{P}\!_{h^{-1}}\hat{f})(\tfrac{k}{p})\,.$ (10)
We may also write the second identity in terms of $(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}$.
Proof 3.2.
The first identity is the Poisson summation formula applied to the $h^{-1}$-periodization of $\hat{f}$. For the second identity, we partition $\mathbb{Z}$ into residue classes modulo $n$ and use $p=hn$. Then
@@
$\hat{f}_{\frac{1}{p},n}\approx(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}\approx\operatorname{\mathcal{F}}\!f_{h,n}\,.$ (11)
More precisely, we split the pointwise error in two components,
$|\frac{1}{\sqrt{p}}\hat{f}(\tfrac{k}{p})-\operatorname{\mathcal{F}}\!f_{h,n}(k)|\leq\frac{1}{\sqrt{p}}|\hat{f}(\tfrac{k}{p})-\mathcal{P}\!_{h^{-1}}\hat{f}(\tfrac{k}{p})|+|\operatorname{\mathcal{F}}\Big{(}(\mathcal{P}\!_{p}f-f)_{h,n}\Big{)}(k)|\,.$ (12)
Taking norms and using the unitarity of $\operatorname{\mathcal{F}}$ we arrive at
$\begin{split}E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|\leq&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|\operatorname{\mathcal{F}}\Big{(}(\mathcal{P}\!_{p}f-f)_{h,n}\Big{)}\|\\ =&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(\mathcal{P}\!_{p}f-f)_{h,n}\|\,.\end{split}$ (13)
On the right hand side of this inequality we have (i) eliminated $\operatorname{\mathcal{F}}$ and (ii) split the error in a time component and a frequency component. Note that the symmetry of this decomposition matches the symmetry of the error $E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}})$, cf. Appendix A. To derive approximation rates, we must ensure that the samples of $f-\mathcal{P}\!_{p}f$ and $\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f}$ decay sufficiently fast. This can be accomplished by specifying decay conditions on $f$ and $\hat{f}$.
3.2 General weights
We quantify the decay of a function by the use of a weight function $v:\mathbb{R}\rightarrow(0,\infty)$. We consider weights $v$ that satisfy
$v\text{ is even: }v(-x)=v(x)\,,$ (14a)
$v\text{ is nondecreasing, i.e., }v(x)\leq v(y),\text{ for }|x|\leq|y|\,,$ (14b)
$v\text{ is submultiplicative, i.e., }v(x+y)\leq v(x)v(y)\text{ for }x,y\in\mathbb{R}\,,$ (14c)
$v\text{ is inverse square-summable, }\sum_{l\in\mathbb{Z}}{|v(l)|^{-2}}<\infty\,.$ (14d)
The standard examples of weights that satisfy conditions (14a) to (14d) are the polynomial weights $v_{\alpha}=(1+|x|)^{\alpha}$ for $\alpha>\frac{1}{2}$, and the (sub-)exponential weights $v_{r,\alpha}=e^{r|x|^{\alpha}}$, for $0<\alpha\leq 1$ and $r>0$.
@@ 3.2 General weights
and the weighted Lebesgue space $L^{2}_{v}:=L^{2}_{v}(\mathbb{R})$ of all functions $f$ such that $\|f\|_{L^{2}_{v}}=\|fv\|_{L^{2}}$ is finite. For a constant weight $v\equiv 1$ we omit the subscript and write $\ell^{p}$ and $L^{p}$.
@@
$\|f\|_{W(\mathcal{C},\ell^{q}_{v})}:=\bigg{(}\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{q}|v(l)|^{q}\bigg{)}^{1/q}$ (15)
is finite, cf. [9, 10, 12].
If $v$ satisfies (14d) then $\ell^{2}_{v}\subseteq\ell^{1}$, and we derive the continuous embeddings $W(\mathcal{C},\ell^{2}_{v})\subseteq W(\mathcal{C},\ell^{1})\subseteq L^{1}(\mathbb{R})$.
In order to bound the right hand side of the error functional $E^{[n]}_{h}(f)$ in (13), we need to estimate the deviation of $f$ and $\hat{f}$ from their respective periodizations. We derive such bounds in terms of the function
$\Phi_{v}(p):=\bigg{(}2\sum_{m=0}^{\infty}|v(pm+\tfrac{p}{2})|^{-2}\bigg{)}^{1/2}\,.$ (16)
Observe that $\lim_{p\to\infty}\Phi_{v}(p)=0$ always holds for a weight $v$ satisfying (3.2). For polynomial and sub-exponential weights we obtain quantitative bounds on the decay of $\Phi_{v}(p)$ in Lemma 3.7 and in Lemma 3.13.
@@
If $f\in W(\mathcal{C},\ell^{2}_{v})$, and $v$ satisfies (3.2) then
$\|f-\mathcal{P}\!_{p}f\|_{L^{2}(-\frac{p}{2},\frac{p}{2})}$ $\leq v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})},$ (i)
$\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ $\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$ (ii)
@@
Proof 3.4.
(i) We bound $(f-\mathcal{P}\!_{p}f)(x)$ by the Cauchy-Schwartz inequality,
$\bigg{|}\sum_{0\neq l\in\mathbb{Z}}f(x+pl)\bigg{|}^{2}\leq\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}\sum_{0\neq l\in\mathbb{Z}}\left|f(x+pl)\right|^{2}|v(x+pl)|^{2}\,.$ (17)
@@
$\begin{split}\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}&\leq\sum_{0\neq m\in\mathbb{Z}}|v(p|m|-\tfrac{p}{2})|^{-2}\\ &\leq 2\sum_{m=1}^{\infty}|v(pm-\tfrac{p}{2})|^{-2}=|\Phi_{v}(p)|^{2}\,.\end{split}$ (18)
@@
After these preparations we can now formulate our main result. This is an error estimate for the approximation of the Fourier transform by the discrete Fourier transform with very general conditions on the decay of the function and its Fourier transform.
@@
Let $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$. If the weights $v,w$ satisfy (3.2), then
$E^{[n]}_{h}(f)\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+w(1)\Phi_{w}(h^{-1})(1+\tfrac{1}{p})^{\frac{1}{2}}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (19)
Note that for $0<h\leq 1\leq p$ the first term does not depend on $h$ and the second term not on $p$. Therefore
$E^{[n]}_{h}(f)\leq 2^{\frac{1}{2}}v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+2^{\frac{1}{2}}w(1)\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (20)
@@
Proof 3.6.
We apply Lemma 3.3 (ii) to both $f$ and $\hat{f}$. Recall the bound $E^{[n]}_{h}(f)\leq\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ in (13). Lemma 3.3 estimates $\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ and yields the first term of the error in (19). Applying Lemma 3.3 to $\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|$ by interchanging the roles of $f$, $p$, and $h$ with $\hat{f}$, $h^{-1}$, and $p^{-1}$, yields the second term. This concludes the proof.
Theorem 3.5 asserts a bound on the error $E^{[n]}_{h}(f)$ in terms of $\Phi_{v}(p)$ and $\Phi_{w}(h^{-1})$. For polynomial and sub-exponential weights the error estimates can be made much more explicit.

### v2

@@ 3.2 Exponential decay in time and polynomial decay in frequency
Figure 6: Logarithmic plots of $E^{[n]}_{h}(f)$ for $f(x)=e^{-2\pi|x|}$. Points are computed numerically for $n=2^{l}$, $l=7,\ldots,23$. We choose $p=n^{\frac{1}{2}}$, $p=6n^{1/10}$, and $p=10$. Reference lines represent the theoretical slope $h^{\beta}$ with $\beta=\frac{3}{2}$ that lead to $n^{-\frac{3}{4}}$, $n^{-\frac{27}{20}}$, and $n^{-\frac{3}{2}}$, respectively. Points align well with theoretical predictions.
This example is also considered in a case study in [5]. For $\delta=0$ the pointwise error is calculated to be of the order $n^{-2}$. However, as soon as $\delta\neq 0$ and $p=10$, say, the required condition $f_{\delta}(p/2)=f_{\delta}(-p/2)$ in [5] is violated, and their bounds yield only $n^{-1}$ for the maximum norm, and hence only $n^{-1/2}$ for the $2$-norm as used in $E^{[n]}_{h}(f)$, whereas both Theorem 2.4 and our numerical experiments consistently yield the error rate $n^{-3/2}$. This further supports the sharpness of our theoretical bounds.
@@ 4 Approximation rates
In this section we prove Theorems 2.1, 2.3, and 2.4 from the introduction. All these statements follow from a single theorem by feeding it with various decay conditions. Specifically, we derive error bounds for the averaged $\ell^{2}$-error $E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|$. This error exhibits a striking symmetry between time and frequency, which will be reflected in the symmetry of the error estimates.
Lemma 4.1.
If $p=nh$ and $f,\hat{f}\in L^{1}(\mathbb{R})$, then
[…]
Proof 4.2.
[…]
The discrete Fourier transform also satisfies $\mathcal{F}^{*}\bar{y}=\overline{\mathcal{F}y}$, where complex conjugation is meant entry-wise. Therefore, we derive $E^{[n]}_{p^{-1}}(\bar{\hat{f}})=\|\overline{\mathcal{F}f_{h,n}}-\bar{\hat{f}}_{\frac{1}{p},n}\|$. Complex conjugation does not affect the norm, so that the latter coincides with $\|\mathcal{F}f_{h,n}-\hat{f}_{\frac{1}{p},n}\|=E^{[n]}_{h}(f)$.
Thus, we expect good bounds for $E^{[n]}_{h}(f)$ to reflect the same symmetry between $(f,h)$ and $(\hat{f},p^{-1})$.
@@ 4.1 Error decomposition into time and frequency components
We derive general and explicit bounds on the approximation error $E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|$ in (6) by decomposing the error into a time and a frequency component that are estimated separately. These components arise from approximating $\hat{f}(\tfrac{k}{p})$ first by a Riemann sum and then truncating it as in (3),
(16) $\hat{f}\big(\tfrac{k}{p}\big)\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}\approx h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=\sqrt{hn}(\operatorname{\mathcal{F}}\!f_{h,n})(k)\,.$
The left-hand side is the sampled (continuous) Fourier transform, the right-hand side is its approximation by the discrete Fourier transform. The idea is to estimate the discretization error
$\hat{f}\big(\tfrac{k}{p}\big)-h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$
by the decay of $\hat{f}$, and the truncation error
$h\sum_{j\in\mathbb{Z}\setminus[n]}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$
by the decay of $f$.
4.2 Wiener amalgam spaces
The treatment of the error makes use of sampling a function on a grid and a strong pointwise version of the Poisson summation formula. To handle these, we introduce the Wiener amalgam spaces that are tailored to these objectives.
The continuous functions on the real line are denoted by $\mathcal{C}=\mathcal{C}(\mathbb{R})$. The space $W(\mathcal{C},\ell^{1})$ consists of all continuous functions $f$ such that
(17) $\|f\|_{W(\mathcal{C},\ell^{1})}:=\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|$
is finite. Wiener amalgam spaces were introduced by N. Wiener [33] and have become convenient and important function spaces in Fourier analysis, cf. [12, 13, 15]. The norm in (17) is easy to understand. A function $f$ belongs to $W(\mathcal{C},\ell^{1})$, if it is continuous and is majorized by an integrable step function with jumps at the integers. See Figure 7.
Figure 7: Visualization of the construction of the Wiener amalgam norm of a continuous function $f$.
In the context of the discretization of the Fourier transform this space arises naturally, because both sampling and periodization are well defined. If $f\in W(\mathcal{C},\ell^{1})$, then the sampling operator $f\mapsto(f(hj))_{j\in\mathbb{Z}}$ maps $W(\mathcal{C},\ell^{1})$ to $\ell^{1}(\mathbb{Z})$ for all $h>0$, and Here $\lceil\cdot\rceil$ denotes the ceiling function.
$\sum_{j\in\mathbb{Z}}|f(jh)|\leq\lceil\tfrac{1}{h}\rceil\|f\|_{W(\mathcal{C},\ell^{1})}\,$,
Consequently the $p$-periodization
@@ 4.2 Wiener amalgam spaces
converges absolutely and uniformly for every $p>0$.
Furthermore, if $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then the Poisson summation formula
$\sum_{l\in\mathbb{Z}}f(x+pl)=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi i\frac{k}{p}x}$
holds pointwise for all $x\in\mathbb{R}$, cf. [18, Lemma 4]. No simple, larger space seems to be known on which the Poisson summation formula holds pointwise for all periods $p$.
@@
(18) $h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}=h\sum_{j\in[n]}(\mathcal{P}\!_{p}f)(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=(\mathcal{P}\!_{h^{-1}}\hat{f})(\tfrac{k}{p})\,.$
We may also write the second identity as $\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}=(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}$.
Proof 4.4.
For the first identity, we partition $\mathbb{Z}$ into residue classes modulo $n$ and use $p=hn$. Then
@@
The second identity in (18) follows from the Poisson summation formula applied to the $h^{-1}$-periodization of $\hat{f}$,
$(\mathcal{P}\!_{h^{-1}}\hat{f})(\tfrac{k}{p})$ $=h\sum_{j\in\mathbb{Z}}\hat{\hat{f}}(hj)e^{2\pi\mathrm{i}hj\frac{k}{p}}$
$=h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}hj\frac{k}{p}}\,.$
This is again the left-hand-side of (18).
@@
(19) $\hat{f}_{\frac{1}{p},n}\approx(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}\approx\operatorname{\mathcal{F}}\!f_{h,n}\,.$
We therefore split the pointwise error in two components,
(20) $|\frac{1}{\sqrt{p}}\hat{f}(\tfrac{k}{p})-\operatorname{\mathcal{F}}\!f_{h,n}(k)|\leq\frac{1}{\sqrt{p}}|\hat{f}(\tfrac{k}{p})-\mathcal{P}\!_{h^{-1}}\hat{f}(\tfrac{k}{p})|+|\operatorname{\mathcal{F}}\big((\mathcal{P}\!_{p}f-f)_{h,n}\big)(k)|\,.$
[…]
(21) $\begin{split}E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|\leq&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|\operatorname{\mathcal{F}}(\mathcal{P}\!_{p}f-f)_{h,n}\|\\ =&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(\mathcal{P}\!_{p}f-f)_{h,n}\|\,.\end{split}$
[…]
To handle the sampling of a function on a grid simultaneously with its decay properties, we introduce another family of Wiener amalgam spaces. We first quantify the decay of a function by the use of a weight function $v:\mathbb{R}\rightarrow(0,\infty)$. We consider weights $v$ that satisfyThe condition (22d) is crucial, whereas the conditions (22a), (22b), and (22c) can be weakened at the expense of more complicated constants in the error estimates.
(22a) v $(x),\displaystyle v\text{ is even: }v$(-x)=v(x)\,,
[…]
(22c) $v\text{ is submultiplicative, i.e., }v(x+y)\leq v(x)v(y)\text{ for }x,y\in\mathbb{R}\,$,
(22d) $v^{-1}\text{ is square-summable, }\sum_{l\in\mathbb{Z}}{|v(l)|^{-2}}<\infty\,.$
The standard examples of weights that satisfy conditions (22a) to (22d) are the polynomial weights $v_{\alpha}(x)=(1+|x|)^{\alpha}$ for $\alpha>\frac{1}{2}$, and the (sub-)exponential weights $v_{r,\alpha}(x)=e^{r|x|^{\alpha}}$, for $0<\alpha\leq 1$ and $r>0$.
@@
and the weighted Lebesgue space $L^{2}_{v}:=L^{2}_{v}(\mathbb{R})$ of all functions $f$ such that $\|f\|_{L^{2}_{v}}=\|fv\|_{L^{2}}$ is finite. For the constant weight $v\equiv 1$ we omit the subscript and write $\ell^{p}$ and $L^{p}$.
@@
(23) $\|f\|_{W(\mathcal{C},\ell^{q}_{v})}:=\bigg(\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{q}|v(l)|^{q}\bigg)^{1/q}$
is finite. For the theory and applications of Wiener amalgam spaces we recommend  [12, 13, 15].
If $v$ satisfies (22d) then $\ell^{2}_{v}\subseteq\ell^{1}$, and we obtain the continuous embeddings $W(\mathcal{C},\ell^{2}_{v})\subseteq W(\mathcal{C},\ell^{1})\subseteq L^{1}(\mathbb{R})$.
4.3 Towards the main error estimates
In order to bound the right-hand side of the error functional $E^{[n]}_{h}(f)$ in (21), we need to estimate the deviation of $f$ and $\hat{f}$ from their respective periodizations. We derive such bounds in terms of the function
(24) $\Phi_{v}(p):=\bigg(2\sum_{m=0}^{\infty}|v(pm+\tfrac{p}{2})|^{-2}\bigg)^{1/2}\,.$
Observe that $\lim_{p\to\infty}\Phi_{v}(p)=0$ always holds for a weight $v$ satisfying (22). For polynomial and sub-exponential weights we will obtain quantitative bounds on the decay of $\Phi_{v}(p)$ in Lemma 4.9 and in Lemma 4.16.
@@ 4.3 Towards the main error estimates
If $f\in W(\mathcal{C},\ell^{2}_{v})$ for a weight $v$ satisfying (22), then
(i) $\|f-\mathcal{P}\!_{p}f\|_{L^{2}(-\frac{p}{2},\frac{p}{2})}$ $\leq v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})},$
(ii) $\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ $\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$
@@
Proof 4.6.
(i) We bound $(f-\mathcal{P}\!_{p}f)(x)=\sum_{0\neq l\in\mathbb{Z}}f(x+pl)$ by the Cauchy-Schwarz inequality,
(25) $\bigg|\sum_{0\neq l\in\mathbb{Z}}f(x+pl)\bigg|^{2}\leq\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}\sum_{0\neq l\in\mathbb{Z}}\left|f(x+pl)\right|^{2}|v(x+pl)|^{2}\,.$
@@
(26) $\begin{split}\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}&\leq\sum_{0\neq m\in\mathbb{Z}}|v(p|m|-\tfrac{p}{2})|^{-2}\\ &\leq 2\sum_{m=1}^{\infty}|v(pm-\tfrac{p}{2})|^{-2}=|\Phi_{v}(p)|^{2}\,.\end{split}$
@@
4.4 The main theorem
After these preparations we can now formulate our main result. This is an error estimate for the approximation of the Fourier transform by the discrete Fourier transform under very general conditions on the decay of the function and its Fourier transform.
@@ 4.4 The main theorem
Let $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$. If the weights $v,w$ satisfy (22), then
(27) $E^{[n]}_{h}(f)\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+w(1)\Phi_{w}(h^{-1})(1+\tfrac{1}{p})^{\frac{1}{2}}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
For $0<h\leq 1\leq p$, we obtain
(28) $E^{[n]}_{h}(f)\leq 2^{\frac{1}{2}}v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+2^{\frac{1}{2}}w(1)\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,$,
so that the first term does not depend on $h$ and the second term not on $p$.
@@
Proof 4.8.
Recall the bound $E^{[n]}_{h}(f)\leq\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ in (21). We apply Lemma 4.5 (ii) to both $f$ and $\hat{f}$. Lemma 4.5 estimates $\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ and yields the first term of the error in (27). Applying Lemma 4.5 to $\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|$ and interchanging the roles of $f$, $p$, and $h$ with $\hat{f}$, $h^{-1}$, and $p^{-1}$, yields the second term. This concludes the proof.
Theorem 4.7 asserts a bound on the error $E^{[n]}_{h}(f)$ in terms of $\Phi_{v}(p)$ and $\Phi_{w}(h^{-1})$. For useful conclusions we need to unravel the error function $\Phi_{v}$, and we will do this for polynomial and sub-exponential weights. The resulting error estimates can then be made completely explicit.